25 ml AgNO3 reacts with 5 mM NaCl; 35 ml of the silver solution is required to titrate a sample of ZnCl2∙6H2O. How many milligrams of the hydrate are present in the sample?
1
Expert's answer
2020-05-13T14:19:33-0400

25 ml AgNO3 reacts with 5 mM (Cl-)

"{\frac {25}{5}}={\frac {35}{x}}"

x = 7 mM.

n(ZnCl2*6H2O) = 7/2 = 3.5 mmol.

m(ZnCl2*6H2O) = 3.5*(65+71+6*18)/1000 = 0.854 g = 854 mg